OpenStudy (anonymous):

Multiply these complex numbers (7-3i) (2-i) Please show the steps

OpenStudy (jdoe0001):

well... what do you get?

OpenStudy (jdoe0001):

is just a pretty straightforward FOIL or just plain multiplication

OpenStudy (anonymous):

i way off, i get 17+i

OpenStudy (jdoe0001):

\(\bf \color{blue}{7-3i}) (2-i)\implies \color{blue}{7}\cdot (2-i)+(\color{blue}{-3i}\cdot (2-i)\)

OpenStudy (jdoe0001):

what does that give you?

OpenStudy (anonymous):

14-i-6i-i ?

OpenStudy (jdoe0001):

hmmm you need to DISTRIBUTE the 7 to both terms inside the parenthesis

OpenStudy (anonymous):

14 -7i -6i^2

OpenStudy (jdoe0001):

\(\bf (7-3i) (2-i)\implies 7\cdot (2-i)+(-3i\cdot (2-i)\\ \quad \\ (7\cdot 2)+(7\cdot -i)+(-3i\cdot 2)+(-3i\cdot -i) \)

OpenStudy (anonymous):

wouldn't i x i equal i^2 ?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

so how would i do that with -i ?

OpenStudy (jdoe0001):

\(\bf (7-3i) (2-i)\implies 7\cdot (2-i)+(-3i\cdot (2-i)\\ \quad \\ (7\cdot 2)+(7\cdot -i)+(-3i\cdot 2)+(-3i\cdot -i)\\ \quad \\ 14-7i-6i+3i^2\\ \quad \\ \textit{recall that}\\ \quad \\ \color{blue}{i^2=\sqrt{-1}\cdot \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies -1}\qquad thus\\ \quad \\ 14-7i-6i+3i^2\implies 14-7i-6i+3(-1)\implies 14-7i-6i-3\\ \quad \\ 11-13i\)

OpenStudy (anonymous):

i got everything up to the -3, for some reason i keep getting positive 3

OpenStudy (jdoe0001):

- times + = - - times - = + same signs multiplication gives POSITIVE different signs multiplication gives NEGATIVE +3 * -1 = -3